# AerialMind: Towards Referring Multi-Object Tracking in UAV Scenarios

### v1

@@ Introduction
Referring Multi-Object Tracking (RMOT) (refer-kitti; refer-kittiv2) aims to achieve precise detection and tracking of specified targets in video sequences through language instructions. It realizes a fundamental paradigm shift from passive perception to active understanding. Although significant progress (refer-dance; refer-bdd; mls; refer-kitti; delving) has been achieved, it is almost entirely confined to ground-level scenarios. It constrains their ability to capture broad-scale scene contexts and perform comprehensive tracking and path planning. In contrast, Unmanned Aerial Vehicles (UAVs) leverage expansive aerial perspectives and superior maneuverability to enable wide-area surveillance capabilities unattainable by ground-based systems. As critical platforms for Embodied AI (wang2025uav), UAVs drive unprecedented demand for intelligent aerial systems with natural language interaction capabilities. However, current RMOT research lacks sufficient exploration of challenging aerial scenarios, resulting in limited real-world applicability and hindering the realization of truly aerial intelligence.
@@
To this end, we construct the first large-scale referring multi-object tracking dataset AerialMind for UAV scenarios. The dataset is extended based on VisDrone (visdrone) and UAVDT (UAVDT), covering multiple flight altitudes, environmental conditions, and target categories. As shown in Figure 1, AerialMind brings unprecedented challenges: ❶ Drastic Appearance Difference: Changes in flight altitude and viewpoints cause dramatic differences in object appearance; ❷ Complex Spatial Relationships: Object relationships under aerial view perspectives are more intricate; ❸ Dynamic Scene Variations: The high maneuverability of UAVs brings continuously changing scenes and illumination conditions; ❹ Various Referring Expressions: Spatial, motion states, and object descriptions in UAV scenarios exhibit richer semantic complexity. To facilitate profound and quantitative analysis, we also pioneer frame-by-frame attribute annotations in the RMOT field.
@@
Table 1: Comparison of referring multi-object tracking datasets. The Dom. represents cross-domain scenarios, Reas. denotes complex reasoning expressions, and Attr. indicates attribute annotation. LaMOT* (lamot) represents the UAV subset.
@@ 1. Refer to caption
Referring to understanding tasks (ding2022vlt; ding2023mose; ding2025mevis; MOSEv2; guan2024nanomvg; guan2025talk2radar; guan2025yu), which aim to localize specific regions in images or videos through natural language expressions. Early dataset construction work mainly focused on static image scenarios, such as the RefCOCO (refcoco) series datasets. Subsequently, researchers gradually extended referring understanding to temporal video domains, successively proposing video referring segmentation datasets such as Refer-DAVIS17 (refer-davis) and Refer-Youtube-VOS (refer-youtube). Wu et al.  (refer-kitti) first proposed the referring multi-object tracking task. Researchers further extended this work, proposing larger-scale Refer-KITTI-V2 (refer-kittiv2), Refer-BDD (refer-bdd) and ReaMOT (reamot) datasets. They mainly focus on specific ground perspectives, lacking sufficient consideration for the unique challenges of aerial platforms such as UAVs. Recently, Researchers (refdrone; liu2025aerialvg) constructed the UAV referring expression detection datasets, validating the feasibility of referring understanding from aerial perspectives. However, they concentrate on single-frame detection tasks, lacking in-depth exploration that requires long-term temporal modeling and complex language understanding.
@@
Early referring understanding methods (refer-davis; early2) mostly adopted two-stage strategies (early4), which rely heavily on candidate region quality and have low computational efficiency. Currently, end-to-end methods (onestage1; onestage2) have gradually become mainstream. These methods achieve visual-language fusion through designing sophisticated mechanisms (onestage3; onestage4; video1; video2; video3). For referring to multi-object tracking, TransRMOT (refer-kitti) first proposed an end-to-end solution based on the Transformer. TempRMOT (refer-kittiv2) further introduced temporal enhancement modules, improving the temporal consistency of tracking. Although these methods (refer-dance; refer-bdd) have achieved significant progress in ground scenarios, they still show inadequate adaptability when facing unique UAV challenges.
@@
In language-guided visual perception tasks, achieving efficient Cross-Modal Representation Alignment (chen2024sato; chen2025ant; chen2025free) is the core challenge. Existing methods mostly follow the early fusion and late fusion paradigms. Early fusion attempts to forcibly align highly abstract text with unstructured, noisy visual features at the beginning of visual encoding. It faces a huge modality gap and may cause the language signal to be progressively diluted in the subsequent encoding. Conversely, although late fusion structures the structural visual features, it makes this process a blind exploration without language navigation, resulting in the final fusion to an inefficient “post-hoc correction”. These become increasingly prominent when facing various descriptions in AerialMind that are full of complex spatial relationships. To this end, we propose a novel Co-evolutionary Fusion Encoder (CFE). Our key insight is: the structuring process of visual features and the guiding process of language information should not be independent stages, but rather a deeply intertwined and mutually reinforcing unified body.
@@
A severe challenge in UAV visual perception is the performance degradation in detecting small-scale objects. Although the Deformable DETR architecture bypasses the traditional FPN, it has an inherent shortcoming. Specifically, the high-resolution feature maps, which are crucial for localizing small objects, possess a severely limited Effective Receptive Field. This results in a significant deficiency of contextual information, making it difficult for the model to distinguish small objects from complex background noise (song2022improving; wang2024windb). To address this, we insert a lightweight yet efficient module, named the Scale-Adaptive Contextual Refinement (SACR), between the encoder and decoder, as shown in Figure 4-B.
@@
The main architectural settings follow those in (refer-kitti). The entire training is deployed on 8 NVIDIA A100 GPUs with a batch size of 1 for 100 epochs. We filtered the target bounding boxes by applying a score threshold of 0.5 and a referring matching score threshold $\beta_{ref}=0.4$. AerialMind utilizes 63 sequences from the VisDrone for the training set and the remaining 17 sequences for in-domain testing. Additionally, we select 13 representative sequences from the UAVDT to serve as the cross-domain test set.
@@
To evaluate the overall tracking performance on AerialMind, we adopt the standard Higher Order Tracking Accuracy $\text{HOTA}=\sqrt{\text{DetA}\cdot\text{AssA}}$ metric (luiten2021hota). To facilitate deeper and more fine-grained diagnostic analysis of model performance, we introduce two attribute-based composite metrics: $\text{HOTA}_{S}$ (Scene-Robustness) and $\text{HOTA}_{M}$ (Motion-Resilience). For $\text{HOTA}_{S}$, the set of attributes $\{A_{i}\}$ comprises Night, Occlusion, and Low Resolution; and the attributes of $\text{HOTA}_{M}$ comprise Viewpoint Change, Scale Variation, Fast Motion, and Rotation. The general formula of attribute-based metrics is: $\text{HOTA}_{A}=\sqrt[N]{\prod_{i=1}^{N}\text{HOTA}_{A_{i}}}$, $N$ denotes the number of attributes included.
@@
We conduct extensive comparisons with state-of-the-art RMOT methods (MOTRV2 (motrv2), TransRMOT (refer-kitti), TempRMOT (refer-kittiv2), CDRMT (cdrmt), MGLT (refer-bdd)). The detailed results are presented in Table 2.
@@
Ground-level Evaluation. As shown in Table 3, we compare with state-of-the-art methods like SKTrack (refer-sktrack) and HFF-Track (zhao2025hff) on the complex expressions ground-level referring dataset Refer-KITTI-V2. Our method also demonstrates competitive performance (35.40% HOTA). It validates that our method provides universal benefits for referring understanding.
@@ Acknowledgments
Figure 7: Overview of the AerialMind, including complex aerial scenarios, large-scale annotations, and diverse evaluation.
@@ Appendix A Appendix
AerialMind contains 48,485 frames, which has a moderate scale compared to some existing datasets (refer-kitti; refer-bdd). It reflects our deliberate focus on scene diversity rather than temporal redundancy. Unlike Refer-Dance (refer-dance) and CRMOT (refer-cross), which achieve high frame counts through repetitive indoor scenarios (dance studios/stages) or multi-viewpoint captures of limited environments, our dataset encompasses over 70 distinct aerial scenarios, providing richer environmental diversity and visual complexity.
Our dataset achieves 247.4 expressions per sequence, ranking second among all benchmarks, which reflects our emphasis on sophisticated semantic interactions rather than simplistic lexical permutations (refer-kittiv2). Our commitment is to capturing the nuanced linguistic complexity inherent in real-world aerial operations. Most significantly, AerialMind leads in both distinct expressions (7,601) and distinct instances (8,778), demonstrating unprecedented semantic richness and object diversity. This achievement validates our dataset’s capacity to encompass the full spectrum of aerial referring scenarios, from fine-grained object discrimination to complex multi-target reasoning tasks.
The temporal ratio per expression of 0.707 reveals another crucial design principle: our referring events are temporally dynamic rather than persistent throughout entire sequences. This characteristic reflects the realistic nature of aerial surveillance and monitoring tasks, where target objects frequently enter and exit the field of view due to UAV mobility and scene complexity. Unlike datasets with near-complete temporal coverage (e.g., Refer-KITTIV2 (refer-kittiv2) at 0.988), our design captures the intermittent and event-driven nature of real-world aerial operations.
@@
The design of our evaluation metrics framework is grounded in a fundamental understanding of the dual nature of challenges inherent in UAV-based referring multi-object tracking systems. Traditional RMOT (refer-kitti; mls) evaluation relies primarily on the standard HOTA metric (luiten2021hota), which provides a balanced assessment of detection and association capabilities through the geometric mean $\text{HOTA}=\sqrt{\text{DetA}\cdot\text{AssA}}$. While HOTA offers valuable insights into overall system performance, it fails to capture the nuanced challenge taxonomy that distinguishes aerial platforms from ground-based systems. It limits our ability to conduct targeted diagnostic analysis and identify specific failure modes in complex UAV scenarios.
@@
Our model employs ResNet50 (resnet) as the visual backbone and RoBERTa (roberta) as the language encoder. Following established practices in referring multi-object tracking, the multi-scale feature maps extracted by the visual backbone undergo encoder-decoder processing for comprehensive cross-modal fusion and spatial relationship modeling. The remaining architectural configurations adhere to the framework established in (refer-kitti).
The training process spans 100 epochs using the AdamW (adaw1) optimizer with an initial learning rate of $1\times 10^{-4}$. The learning rate undergoes decay by a factor of 10 at the $40^{th}$ epoch to ensure convergence stability. The loss function incorporates multiple weighted components: $\lambda_{cls}$, $\lambda_{L_{1}}$, $\lambda_{giou}$, and $\lambda_{align}$ are configured to 2, 5, 2, and 2, respectively, to balance classification accuracy, localization precision, and alignment quality. To accommodate the diverse linguistic expressions characteristic of aerial scenarios, we initialize the framework with 300 object queries. To ensure reproducibility and fair comparison across different runs, we maintain a fixed random seed throughout both training and testing phases. Our HETrack model with 51.4M trainable parameters achieves 15.6 FPS on a single RTX 4080 during inference.
@@
Co-evolutionary Fusion Encoder: Information-Theoretic Motivation. Traditional fusion paradigms suffer from the information bottleneck problem (lai2023faithful; lai2025learning), where cross-modal alignment is constrained by unidirectional information flow. Let $V$ and $L$ represent visual and linguistic feature distributions, respectively. Conventional approaches are limited by $I(V;L)\leq\min(H(V),H(L))$, where mutual information is bounded by the entropy of individual modalities. Our CFE addresses this fundamental limitation through bidirectional iterative. It enables visual and linguistic representations to evolve synergistically rather than independently.
Scale Adaptive Contextual Refinement: Multi-scale Information Theory. The SACR module is motivated by multi-scale information processing theory (lindeberg2013scale), which posits that optimal feature representation for small object detection requires principled information integration across multiple spatial scales. For UAV scenarios with extreme scale variations, the effective receptive ($ERF_{optimal}$) field must satisfy:
@@
where $F_{r}^{(s)}$ represents features at scale $s$ with receptive field $r$, and $MI$ denotes mutual information with target labels $Y$. Our atrous convolution configuration with dilation rates $\{6,12,18\}$ approximates this optimal multi-scale integration while preserving spatial resolution. The adaptive channel recalibration mechanism seeks to emphasize task-relevant features while reducing noise interference (guan2025watervg) for robust representation.
@@
To further validate the generalizability and robustness of our proposed HETrack method, we conduct comprehensive evaluations on the Refer-KITTI-V2 dataset (refer-kittiv2). It represents the most extensive ground-perspective RMOT benchmark featuring complex linguistic descriptions. It’s a crucial evaluation for our method’s adaptability across diverse scenarios and semantic complexities.
@@
To provide comprehensive insights into the practical effectiveness of our proposed HETrack method, we present detailed qualitative comparisons with state-of-the-art approaches TempRMOT (refer-kittiv2) and MGLT (refer-bdd) across challenging scenarios in Figure 9-11.
@@
As shown in Figure 9, the basketball court scene presents a particularly demanding test case involving dense crowds, frequent occlusions, and rapid multi-directional movements. Both TempRMOT (refer-kittiv2) and MGLT (refer-bdd) exhibit failures in detecting “individuals wearing dark upper garments” in this complex scene. In stark contrast, HETrack robustly detects and tracks all relevant individuals.
@@
The nighttime intersection scene presents a more sophisticated challenge, requiring accurate interpretation of motion states rather than visual appearance matching. The referring expression “white automobiles in stationary condition” demands precise discrimination between stationary and moving vehicles of the same color category. TempRMOT (refer-kittiv2) demonstrates semantic confusion by incorrectly identifying moving white vehicles as targets, while MGLT (refer-bdd) suffers from incomplete detection, missing several legitimate stationary targets. HETrack achieves superior semantic precision by correctly identifying only the stationary white automobiles.

### v2

@@ Introduction
Referring Multi-Object Tracking (RMOT) (Wu et al. 2023; Zhang et al. 2024) aims to achieve precise detection and tracking of specified targets in video sequences through language instructions. It realizes a fundamental paradigm shift from passive perception to active understanding. Although significant progress (Du et al. 2024; Chen et al. 2025a; Ma et al. 2024; Wu et al. 2023; Chen et al. 2024a) has been achieved, it is almost entirely confined to ground-level scenarios. It constrains their ability to capture broad-scale scene contexts and perform comprehensive tracking and path planning. In contrast, Unmanned Aerial Vehicles (UAVs) leverage expansive aerial perspectives and superior maneuverability to enable wide-area surveillance capabilities unattainable by ground-based systems. As critical platforms for Embodied AI (Wang et al. 2025), UAVs drive unprecedented demand for intelligent aerial systems with natural language interaction capabilities. However, current RMOT research lacks sufficient exploration of challenging aerial scenarios, resulting in limited real-world applicability and hindering the realization of truly aerial intelligence.
@@
To this end, we construct the first large-scale referring multi-object tracking dataset AerialMind for UAV scenarios. The dataset is extended based on VisDrone (Du et al. 2019) and UAVDT (Du et al. 2018), covering multiple flight altitudes, environmental conditions, and target categories. As shown in Figure 1, AerialMind brings unprecedented challenges: ❶ Drastic Appearance Difference: Changes in flight altitude and viewpoints cause dramatic differences in object appearance; ❷ Complex Spatial Relationships: Object relationships under aerial view perspectives are more intricate; ❸ Dynamic Scene Variations: The high maneuverability of UAVs brings continuously changing scenes and illumination conditions; ❹ Various Referring Expressions: Spatial, motion states, and object descriptions in UAV scenarios exhibit richer semantic complexity. To facilitate profound and quantitative analysis, we also pioneer frame-by-frame attribute annotations in the RMOT field.
@@
Table 1: Comparison of referring multi-object tracking datasets. The Dom. represents cross-domain scenarios, Reas. denotes complex reasoning expressions, and Attr. indicates attribute annotation. LaMOT* (Li et al. 2025a) represents the UAV subset.
@@ 1. Refer to caption
Referring to understanding tasks (Ding et al. 2022a, 2023, 2025a, 2025b; Guan et al. 2024, 2025c, 2025b), which aim to localize specific regions in images or videos through natural language expressions. Early dataset construction work mainly focused on static image scenarios, such as the RefCOCO (Yu et al. 2016) series datasets. Subsequently, researchers gradually extended referring understanding to temporal video domains, successively proposing video referring segmentation datasets such as Refer-DAVIS17 (Khoreva, Rohrbach, and Schiele 2019) and Refer-Youtube-VOS (Seo, Lee, and Han 2020). Wu et al.  (Wu et al. 2023) first proposed the referring multi-object tracking task. Researchers further extended this work, proposing larger-scale Refer-KITTI-V2 (Zhang et al. 2024), Refer-BDD (Chen et al. 2025a) and ReaMOT (Chen et al. 2025b) datasets. They mainly focus on specific ground perspectives, lacking sufficient consideration for the unique challenges of aerial platforms such as UAVs. Recently, Researchers (Sun et al. 2025; Liu et al. 2025) constructed the UAV referring expression detection datasets, validating the feasibility of referring understanding from aerial perspectives. However, they concentrate on single-frame detection tasks, lacking in-depth exploration that requires long-term temporal modeling and complex language understanding.
@@
Early referring understanding methods (Khoreva, Rohrbach, and Schiele 2019; Luo and Shakhnarovich 2017) mostly adopted two-stage strategies (Zhou et al. 2022), which rely heavily on candidate region quality and have low computational efficiency. Currently, end-to-end methods (Liang et al. 2023; Liao et al. 2020) have gradually become mainstream. These methods achieve visual-language fusion through designing sophisticated mechanisms (Luo et al. 2020; Sun et al. 2020; Ding et al. 2022b; Hui et al. 2021; Wu et al. 2022). For referring to multi-object tracking, TransRMOT (Wu et al. 2023) first proposed an end-to-end solution based on the Transformer. TempRMOT (Zhang et al. 2024) further introduced temporal enhancement modules, improving the temporal consistency of tracking. Although these methods (Du et al. 2024; Chen et al. 2025a) have achieved significant progress in ground scenarios, they still show inadequate adaptability when facing unique UAV challenges.
@@
In language-guided visual perception tasks, achieving efficient Cross-Modal Representation Alignment (Chen et al. 2024b, 2025d, 2025c) is the core challenge. Existing methods mostly follow the early fusion and late fusion paradigms. Early fusion attempts to forcibly align highly abstract text with unstructured, noisy visual features at the beginning of visual encoding. It faces a huge modality gap and may cause the language signal to be progressively diluted in the subsequent encoding. Conversely, although late fusion structures the structural visual features, it makes this process a blind exploration without language navigation, resulting in the final fusion to an inefficient “post-hoc correction”. These become increasingly prominent when facing various descriptions in AerialMind that are full of complex spatial relationships. To this end, we propose a novel Co-evolutionary Fusion Encoder (CFE). Our key insight is: the structuring process of visual features and the guiding process of language information should not be independent stages, but rather a deeply intertwined and mutually reinforcing unified body.
@@
A severe challenge in UAV visual perception is the performance degradation in detecting small-scale objects. Although the Deformable DETR architecture bypasses the traditional FPN, it has an inherent shortcoming. Specifically, the high-resolution feature maps, which are crucial for localizing small objects, possess a severely limited Effective Receptive Field. This results in a significant deficiency of contextual information, making it difficult for the model to distinguish small objects from complex background noise (Song et al. 2022; Wang et al. 2024). To address this, we insert a lightweight yet efficient module, named the Scale-Adaptive Contextual Refinement (SACR), between the encoder and decoder, as shown in Figure 4-B.
@@
The main architectural settings follow those in (Wu et al. 2023). The entire training is deployed on 8 NVIDIA A100 GPUs with a batch size of 1 for 100 epochs. We filtered the target bounding boxes by applying a score threshold of 0.5 and a referring matching score threshold $\beta_{ref}=0.4$. AerialMind utilizes 63 sequences from the VisDrone for the training set and the remaining 17 sequences for in-domain testing. Additionally, we select 13 representative sequences from the UAVDT to serve as the cross-domain test set.
@@
To evaluate the overall tracking performance on AerialMind, we adopt the standard Higher Order Tracking Accuracy $\text{HOTA}=\sqrt{\text{DetA}\cdot\text{AssA}}$ metric (Luiten et al. 2021). To facilitate deeper and more fine-grained diagnostic analysis of model performance, we introduce two attribute-based composite metrics: $\text{HOTA}_{S}$ (Scene-Robustness) and $\text{HOTA}_{M}$ (Motion-Resilience). For $\text{HOTA}_{S}$, the set of attributes $\{A_{i}\}$ comprises Night, Occlusion, and Low Resolution; and the attributes of $\text{HOTA}_{M}$ comprise Viewpoint Change, Scale Variation, Fast Motion, and Rotation. The general formula of attribute-based metrics is: $\text{HOTA}_{A}=\sqrt[N]{\prod_{i=1}^{N}\text{HOTA}_{A_{i}}}$, $N$ denotes the number of attributes included.
@@
We conduct extensive comparisons with state-of-the-art RMOT methods (MOTRV2 (Zhang, Wang, and Zhang 2023), TransRMOT (Wu et al. 2023), TempRMOT (Zhang et al. 2024), CDRMT (Liang et al. 2025), MGLT (Chen et al. 2025a)). The detailed results are presented in Table 2.
@@
Ground-level Evaluation. As shown in Table 3, we compare with state-of-the-art methods like SKTrack (Li et al. 2025b) and HFF-Track (Zhao et al. 2025) on the complex expressions ground-level referring dataset Refer-KITTI-V2. Our method also demonstrates competitive performance (35.40% HOTA). It validates that our method provides universal benefits for referring understanding.
@@ Appendix A Appendix
AerialMind contains 48,485 frames, which has a moderate scale compared to some existing datasets (Wu et al. 2023; Chen et al. 2025a). It reflects our deliberate focus on scene diversity rather than temporal redundancy. Unlike Refer-Dance (Du et al. 2024) and CRMOT (Chen, Yu, and Tao 2025), which achieve high frame counts through repetitive indoor scenarios (dance studios/stages) or multi-viewpoint captures of limited environments, our dataset encompasses over 70 distinct aerial scenarios, providing richer environmental diversity and visual complexity.
Our dataset achieves 247.4 expressions per sequence, ranking second among all benchmarks, which reflects our emphasis on sophisticated semantic interactions rather than simplistic lexical permutations (Zhang et al. 2024). Our commitment is to capturing the nuanced linguistic complexity inherent in real-world aerial operations. Most significantly, AerialMind leads in both distinct expressions (7,601) and distinct instances (8,778), demonstrating unprecedented semantic richness and object diversity. This achievement validates our dataset’s capacity to encompass the full spectrum of aerial referring scenarios, from fine-grained object discrimination to complex multi-target reasoning tasks.
The temporal ratio per expression of 0.707 reveals another crucial design principle: our referring events are temporally dynamic rather than persistent throughout entire sequences. This characteristic reflects the realistic nature of aerial surveillance and monitoring tasks, where target objects frequently enter and exit the field of view due to UAV mobility and scene complexity. Unlike datasets with near-complete temporal coverage (e.g., Refer-KITTIV2 (Zhang et al. 2024) at 0.988), our design captures the intermittent and event-driven nature of real-world aerial operations.
@@
The design of our evaluation metrics framework is grounded in a fundamental understanding of the dual nature of challenges inherent in UAV-based referring multi-object tracking systems. Traditional RMOT (Wu et al. 2023; Ma et al. 2024) evaluation relies primarily on the standard HOTA metric (Luiten et al. 2021), which provides a balanced assessment of detection and association capabilities through the geometric mean $\text{HOTA}=\sqrt{\text{DetA}\cdot\text{AssA}}$. While HOTA offers valuable insights into overall system performance, it fails to capture the nuanced challenge taxonomy that distinguishes aerial platforms from ground-based systems. It limits our ability to conduct targeted diagnostic analysis and identify specific failure modes in complex UAV scenarios.
@@
Our model employs ResNet50 (He et al. 2016) as the visual backbone and RoBERTa (Liu et al. 2019) as the language encoder. Following established practices in referring multi-object tracking, the multi-scale feature maps extracted by the visual backbone undergo encoder-decoder processing for comprehensive cross-modal fusion and spatial relationship modeling. The remaining architectural configurations adhere to the framework established in (Wu et al. 2023).
The training process spans 100 epochs using the AdamW (Loshchilov and Hutter 2019) optimizer with an initial learning rate of $1\times 10^{-4}$. The learning rate undergoes decay by a factor of 10 at the $40^{th}$ epoch to ensure convergence stability. The loss function incorporates multiple weighted components: $\lambda_{cls}$, $\lambda_{L_{1}}$, $\lambda_{giou}$, and $\lambda_{align}$ are configured to 2, 5, 2, and 2, respectively, to balance classification accuracy, localization precision, and alignment quality. To accommodate the diverse linguistic expressions characteristic of aerial scenarios, we initialize the framework with 300 object queries. To ensure reproducibility and fair comparison across different runs, we maintain a fixed random seed throughout both training and testing phases. Our HETrack model with 51.4M trainable parameters achieves 15.6 FPS on a single RTX 4080 during inference.
@@
Co-evolutionary Fusion Encoder: Information-Theoretic Motivation. Traditional fusion paradigms suffer from the information bottleneck problem (Lai et al. 2023, 2025), where cross-modal alignment is constrained by unidirectional information flow. Let $V$ and $L$ represent visual and linguistic feature distributions, respectively. Conventional approaches are limited by $I(V;L)\leq\min(H(V),H(L))$, where mutual information is bounded by the entropy of individual modalities. Our CFE addresses this fundamental limitation through bidirectional iteration. It enables visual and linguistic representations to evolve synergistically rather than independently.
Scale Adaptive Contextual Refinement: Multi-scale Information Theory. The SACR module is motivated by multi-scale information processing theory (Lindeberg 2013), which posits that optimal feature representation for small object detection requires principled information integration across multiple spatial scales. For UAV scenarios with extreme scale variations, the effective receptive ($ERF_{optimal}$) field must satisfy:
@@
where $F_{r}^{(s)}$ represents features at scale $s$ with receptive field $r$, and $MI$ denotes mutual information with target labels $Y$. Our atrous convolution configuration with dilation rates $\{6,12,18\}$ approximates this optimal multi-scale integration while preserving spatial resolution. The adaptive channel recalibration mechanism seeks to emphasize task-relevant features while reducing noise interference (Guan et al. 2025a) for robust representation.
@@
To further validate the generalizability and robustness of our proposed HETrack method, we conduct comprehensive evaluations on the Refer-KITTI-V2 dataset (Zhang et al. 2024). It represents the most extensive ground-perspective RMOT benchmark featuring complex linguistic descriptions. It’s a crucial evaluation for our method’s adaptability across diverse scenarios and semantic complexities.
@@
To provide comprehensive insights into the practical effectiveness of our proposed HETrack method, we present detailed qualitative comparisons with state-of-the-art approaches TempRMOT (Zhang et al. 2024) and MGLT (Chen et al. 2025a) across challenging scenarios in Figure 9-11.
@@
As shown in Figure 9, the basketball court scene presents a particularly demanding test case involving dense crowds, frequent occlusions, and rapid multi-directional movements. Both TempRMOT (Zhang et al. 2024) and MGLT (Chen et al. 2025a) exhibit failures in detecting “individuals wearing dark upper garments” in this complex scene. In stark contrast, HETrack robustly detects and tracks all relevant individuals.
@@
The nighttime intersection scene presents a more sophisticated challenge, requiring accurate interpretation of motion states rather than visual appearance matching. The referring expression “white automobiles in stationary condition” demands precise discrimination between stationary and moving vehicles of the same color category. TempRMOT (Zhang et al. 2024) demonstrates semantic confusion by incorrectly identifying moving white vehicles as targets, while MGLT (Chen et al. 2025a) suffers from incomplete detection, missing several legitimate stationary targets. HETrack achieves superior semantic precision by correctly identifying only the stationary white automobiles.